# LiDAR 2.0: Hierarchical Curvy Waveguide Detailed Routing for Large-Scale Photonic Integrated Circuits

### v1

@@ Abstract
Driven by innovations in photonic computing and interconnects, photonic integrated circuit (PIC) designs advance and grow in complexity. Traditional manual physical design processes have become increasingly cumbersome. Available PIC layout tools are mostly schematic-driven, which has not alleviated the burden of manual waveguide planning and layout drawing. Previous research in PIC automated routing are largely adapted from electronic design, focusing on high-level planning and overlooking photonic-specific constraints such as curvy waveguides, bending, and port alignment. As a result, they fail to scale and cannot generate DRV-free layouts, highlighting the need for dedicated electronic-photonic design automation tools to streamline PIC physical design. In this work, we present LiDAR, the first automated PIC detailed router for large-scale designs. It features a grid-based, curvy-aware A∗ engine with adaptive crossing insertion, congestion-aware net ordering, and insertion-loss optimization. To enable routing in more compact and complex designs, we further extend our router to hierarchical routing as LiDAR 2.0. It introduces redundant-bend elimination, crossing space preservation, and routing order refinement for improved conflict resilience. We also develop and open-source a YAML-based PIC intermediate representation and diverse benchmarks, including TeMPO, GWOR, and Bennes, which feature hierarchical structures and high crossing densities. Evaluations across various benchmarks show that LiDAR 2.0 consistently produces DRV-free layouts, achieving up to 16% lower insertion loss and 7.69× speedup over prior methods on spacious cases, and 9% lower insertion loss with 6.95× speedup over LiDAR 1.0 on compact cases. Our codes are open-sourced at link.
@@ I Introduction
In recent years, photonic integrated circuits have attracted considerable attention due to their advantages in high-speed data transmission and low power consumption. Notable advances include photonic tensor cores (PTCs) [1, 2, 3] developed for optical neural networks (ONNs) [3], as well as photonic network-on-chips (NoCs) [4, 5, 6, 7, 8] designed for high-bandwidth on-chip communication. Driven by these emerging applications, PIC designs are rapidly increasing in complexity. As shown in Fig. 1, the number of photonic components per chip is approaching the scale of 1000 and is expected to double approximately every 2.5 years [9]. This exponential growth underscores the urgent need for advanced electronic-photonic design automation (EPDA) toolchains [10, 11, 12] that can streamline layout generation, enhance productivity, and ensure design quality.
@@
Most existing studies primarily focus on global routing planning for PICs. A number of optical routing algorithms have been proposed for on-chip 3D system-on-package designs [16, 17, 18], with the primary goal of optimizing signal loss and overall power consumption. Tools such as PROTON [19] and PLATON [20] offer automatic placement and routing, employing a modified Lee’s algorithm for waveguide routing. In [21], further improvements are achieved by optimizing device orientation, such as flipping and rotation, to reduce insertion loss through fewer crossings. While these global routing techniques are effective in identifying low-loss paths, they often focus solely on logical path planning and tend to neglect physical realizability. As a result, they may suffer from routing congestion, infeasible bend or crossing insertions, ultimately leading to an invalid routing solution.
@@
Extensive evaluations show that LiDAR 2.0 supports PIC routing with various crossing sizes and bend radii, achieving DRV-free layouts with up to 16% lower insertion loss (IL) and 7.69 speedup over prior methods on spacious benchmarks, and nearly DRV-free layouts with 9% lower IL and 6.95 speedup over LiDAR 1.0 on compact benchmarks.
@@ II-B Photonic Design Rules
PIC routing typically operates on a single waveguide layer, where all nets are 2-pin point-to-point optical paths. Below, we summarize key design rules and highlight photonic-specific considerations.
II-B1 Waveguide Spacing
Waveguides must maintain sufficient spacing from each other and from photonic device structures to prevent unwanted crosstalk. As shown in Fig. 2(a), the minimum required spacing depends on various factors such as wavelength, polarization mode, refractive index contrast, substrate type, and waveguide cross-section. For high-index contrast systems (e.g., silicon-on-insulator), relatively small spacings of 1–3 are typically sufficient.
@@ II-B5 Signal Integrity
Insertion loss is a critical metric for evaluating PIC routing quality, as it directly affects the laser power budget and signal integrity, including signal-to-noise ratio and crosstalk as illustrated in Fig. 2(f) The primary evaluation metric is the maximum insertion loss along the critical path. Long waveguides and excessive crossings degrade signal integrity and should be minimized wherever possible.
@@ II-C Schematic-Driven PIC Layout
Traditional PIC layout workflows, including manual design and current EPDA tools, are schematic-driven [37]. In this approach, all structures, including crossings and waveguide segments are instantiated explicitly. Designers must plan routing during the schematic stage and manually insert crossings, while waveguide connections rely on port abutment without explicit physical net definitions.
@@ III LiDAR: Automated PIC Detailed Routing
With an input circuit netlist, as shown in Fig. 3, we present LiDAR, a detailed PIC routing framework based on a customized grid-based A∗ search. It efficiently generates curvy waveguide paths with automatic crossings insertion, aiming to minimize the maximum insertion loss, while ensuring design rules compliance. The core routing procedure consists of the following three main phases: ➊ Port Access Assignment: assigns access ports based on port orientation and local density to facilitate smooth routing and reduce congestion; ➋ Iterative Curvy-Aware Waveguide Routing: route all nets using a curvy-aware A∗ search algorithm, guided by group-based net ordering. ➌ Crossing Optimization and Route Refinement: a local rip-up and reroute mechanism is applied to resolve conflicts and optimize crossings, followed by refinement of the routing solution to eliminate redundant bendings and produce a design-rule-clean GDS layout.
@@ III-A Accessibility-Enhanced Port Assignment
Port Propagation. Some waveguide ports are embedded inside the bounding box of a device. Since devices are treated as routing blockages, we propagate these internal ports outward to the device boundary based on their orientations, as illustrated in Fig. 4(a). This ensures they are reachable during routing.
Bending-Aware Port Access Region Reservation. To prevent other waveguides from blocking access to ports, a region in front of each port is reserved along its orientation as shown in Fig. 4(a). The reserved size is adaptively determined based on the waveguide’s bending radius, ensuring sufficient space for bends and improving the success rate of port access while minimizing area overhead.
Congested Port Spreading. Certain PIC devices feature dense clusters of ports that may overlap on the same routing grid, leading to access conflicts. To alleviate this, we apply a symmetric spreading strategy, where high-density ports are redistributed with a predefined spacing and extension length, as shown in Fig. 4(b). These new access ports are connected to their original locations via sine bends, ensuring they occupy distinct routing tracks and reducing congestion. The reserved access regions are also updated accordingly.
Channel Planning via Staggered Access Point Offsets. To further improve access in dense layouts, we introduce staggered access point regions, particularly beneficial for devices like multimode interference (MMI) structures that have many ports aligned on one side, as shown in Fig. 4(c). High port density may cause access ports to be blocked by nearby waveguides, while closely spaced parallel waveguides can prevent other nets from inserting crossings over them. We address this by progressively extending the access region length for inner ports and introducing an access offset larger than the crossing footprint. This approach provides sufficient space for bends and enables consecutive waveguide crossings, allowing other nets to traverse through parallel routing paths.
@@ III-B Port-Group-based Net Order
III-C Non-Manhattan Waveguide Routing with Curvy-Aware A∗ Search
@@ III-C2 Parametric Curvy-Aware Neighbor Candidate Generation
The ceiling function guarantees that there is enough spacing to accommodate bending structures. Note that unlike 45∘ diagonal neighbors in traditional 8-way A∗ search, where is equal to , our neighbors are the endpoint of the 45-degree curves, where and are not equal.
@@ III-C4 Insertion Loss-Aware A∗ Search Cost
LiDAR adopts a customized A∗ search cost function that incorporates insertion loss considerations to improve both routing quality and algorithm efficiency. The total cost of a node in A∗ search is given by the standard formulation: , where denotes the accumulated cost from the source node to the current node , and is the heuristic estimate of the cost from node to the target node . The cost consists of two components: the insertion loss along the path to the current node, denoted as and calculated based on Eq. (1), and a group-based congestion penalty (GCP), denoted as to save routing resources for a group of nets:
@@
Since our method allows non-Manhattan routing, the shortest path is often diagonal. To ensure the A∗ heuristic remains valid and does not overestimate the cost, we define a customized heuristic as shown in Eq. (5):
where and represent the minimum and maximum differences in the x and y coordinates between the current node and the target node , respectively. Additionally, a penalty term proportional to the 45∘ bending loss, , is added when both and are non-zero, as this scenario requires at least one 45∘ bend. This penalty encourages paths that end in orientations better suited for port connection.
@@ IV-A Hierarchical Netlist Tree Construction
PIC netlists often exhibit inherent hierarchical structure, where complex systems are composed of repeated functional modules as shown in Fig. 11. Leveraging this structure is crucial for scalable physical design. To exploit such hierarchy, we first construct a Hierarchical Tree based on the invocation relationships between modules and the overall circuit. This hierarchical representation enables global routing planning at multiple levels. We then perform routing in a bottom-up manner, completing routing for each level of modules progressively. This multi-level routing approach helps avoid conflicts between different modules, as routing is localized within each submodule before higher-level integration. Similar to standard-cell design in VLSI, the internal routing resources of a module are treated as private and should not be accessed by external nets.
@@ IV-B Routing Conflict Reduction via Spatial-Aware Optimization
Offset Neighbor. Although the algorithm naturally discourages excessive bending through its bend cost heuristic, it may still introduce unnecessary bends when ports are slightly misaligned. Especially in tight spaces, such misalignment can trigger multiple bends, increasing insertion loss and risking routing failure, as shown in Fig. 12(a).
To address this issue, we incorporate additional offset neighbors to avoid such detours under specific conditions: when the current node is close to the target node (within four times the bending radius), and the orientation differs by 180 degrees from the target node. These offset neighbors are designed to correct small misalignments (less than the bending radius) between the source and target ports. The offset neighbor locations are computed analytically based on the predefined bending radius and grid size, ensuring that the resulting path remains smooth and physically realizable using sine-bend waveguides. From an algorithmic perspective, the introduction of offset neighbors simplifies port access, reduces redundant neighbor exploration caused by bending detours, and thereby improves the search efficiency.
@@
where is a user-defined penalty coefficient and denotes the minimum width required to insert a valid crossing structure. The penalty is applied over a -wide region around the net segment, discouraging other nets from occupying this reserved space.
@@ IV-C Reduce Routing Conflict via Routing Order Optimization
Although the port region allows access points to be staggered, reducing conflicts during access, an inappropriate access order may easily result in the target ports being blocked by routed waveguides and other port regions. Especially for multi-port components, the intra-group routing order is critically important. As shown in Fig. 13(a), both the orientation and spatial location of ports significantly influence the routing order, as they jointly determine the accessibility and potential for routing conflicts among nets.
To address the intra-group routing order, we determine the routing order by the spatial relationship between source ports and target ports within the group, as shown in Fig. 13(b). Let the set of source ports be denoted as . We compute their geometric mean coordinates. If each port has a position , the average position is:
After obtaining the mean location of the source ports, we determine the routing order based on the Manhattan distance from each target port to . Target ports that are closer to the average source location are assigned higher routing priority. This approach can effectively resolve most routing conflict issues, particularly in multiport components.
Routing Order Refinement. It is difficult to guarantee an optimal routing order. In particular, when nets are routed independently without group information, the aforementioned method becomes inapplicable. Hence, it is necessary to dynamically adjust the order during the routing process. An inappropriate order may lead to ports being blocked by previously routed waveguides, resulting in escape or access failures. As illustrated in Fig. 14, since net1, net2, and net3 are routed independently, net2 fails because it cannot access the port due to net1’s port region and net3’s waveguide blocking its path.
To address routing failures, we enhance the rip-up and re-route scheme by not only removing nets that directly conflict with the failed path but also ripping up those located in front of the failed net’s ports to alleviate congestion. As shown in Fig. 14, we scan along the port orientation and identify the conflict order and names of the conflicting nets (starting from ), which serve as guidance to adjust the routing order. Higher-priority nets are routed first to improve escape success, but may block others’ access. To avoid this, we dynamically extend the target port regions of high-priority nets, shifting their access points and reducing the risk of subsequent blocking.
Our extended conflict-resilient hierarchical routing scheme can effectively enhance routability even under tight layout constraints and help accelerate the routing process.
@@ V PIC Routing Benchmark
Bennes [45]: Based on Mach-Zehnder interferometers, it adopts a structured, hierarchical Benes topology. It exhibits extremely high crossing density, where a single net may require tens of crossings, challenging both crossing reservation and congestion-aware routing strategies.
Tables II and III summarize the benchmark statistics. Total #CR refers to the number of topological crossings inherent to the circuit and resolved during routing. For Light and GWOR, internal switch crossings are excluded, as we focus solely on the routing between the switch and the I/O interfaces in these benchmarks. Table II (ISPD version) includes spacious, moderately complex layouts, while Table III features more compact benchmarks with higher component density, smaller die sizes, and more crossings, posing greater routing challenges and better reflecting real-world photonic layout constraints. The bending radius is set to 60 for Light and GWOR and 5 for others.

### v2

@@ Abstract
Driven by innovations in photonic computing and interconnects, photonic integrated circuit (PIC) designs advance and grow in complexity. Traditional manual physical design processes have become increasingly cumbersome. Available PIC layout tools are mostly schematic-driven, which has not alleviated the burden of manual waveguide planning and layout drawing. Previous research in PIC automated routing is largely adapted from electronic design, focusing on high-level planning and overlooking photonic-specific constraints such as curvy waveguides, bending, and port alignment. As a result, they fail to scale and cannot generate DRV-free layouts, highlighting the need for dedicated electronic-photonic design automation tools to streamline PIC physical design. In this work, we present LiDAR, the first automated PIC detailed router for large-scale designs. It features a grid-based, curvy-aware A∗ engine with adaptive crossing insertion, congestion-aware net ordering, and insertion-loss optimization. To enable routing in more compact and complex designs, we further extend our router to hierarchical routing as LiDAR 2.0. It introduces redundant-bend elimination, crossing space preservation, and routing order refinement for improved conflict resilience. We also develop and open-source a YAML-based PIC intermediate representation and diverse benchmarks, including TeMPO, GWOR, and Bennes, which feature hierarchical structures and high crossing densities. Evaluations across various benchmarks show that LiDAR 2.0 produces nearly DRV-free layouts, achieving up to 16% lower insertion loss and 7.69× speedup over prior methods on spacious cases, and 9% lower insertion loss with 6.95× speedup over LiDAR 1.0 on compact cases. Our codes are open-sourced at link.
@@ I Introduction
In recent years, photonic integrated circuits have attracted considerable attention due to their advantages in high-speed data transmission and low power consumption. Notable advances include photonic tensor cores (PTCs) [1] developed for optical neural networks (ONNs) [2], as well as photonic network-on-chips (NoCs) [3] designed for high-bandwidth on-chip communication. Driven by these emerging applications, PIC designs are rapidly increasing in complexity. As shown in Fig. 1, the number of photonic components per chip is approaching the scale of 1000 and is expected to double approximately every 2.5 years [4]. This growth highlights the urgent need for advanced electronic-photonic design automation (EPDA) toolchains [5] to streamline layout, boost productivity, and ensure quality.
@@
Most existing studies primarily focus on global routing planning for PICs. A number of optical routing algorithms have been proposed for on-chip 3D system-on-package designs [10, 11], with the primary goal of optimizing signal loss and overall power consumption. Tools such as PROTON [12] and PLATON [13] offer automatic placement and routing, employing a modified Lee’s algorithm for waveguide routing. In [14], further improvements are achieved by optimizing device orientation, such as flipping and rotation, to reduce insertion loss through fewer crossings. While these global routing techniques are effective in identifying low-loss paths, they often focus solely on logical path planning and tend to neglect physical realizability. As a result, they may suffer from routing congestion, infeasible bend or crossing insertions, ultimately leading to an invalid routing solution.
@@
Evaluations show that LiDAR 2.0 supports PIC routing with various crossing sizes and bend radii, achieving DRV-free layouts with 16% lower insertion loss (IL) and 7.69 speedup over prior methods on spacious benchmarks, and nearly DRV-free layouts with 9% lower IL and 6.95 speedup over LiDAR 1.0 on compact benchmarks.
@@ II-B Photonic Design Rules
PIC routing typically operates on a single waveguide layer, where all nets are 2-pin optical paths. We summarize key design rules and highlight photonic-specific considerations.
II-B1 Waveguide Type
PICs often contain multiple types of waveguides, whose characteristics depend on factors such as wavelength, polarization mode, substrate type, and cross-section, as shown in Fig. 2(a). Waveguide connections must be strictly matched in type or transitioned via tapers to minimize loss. Different waveguide types also have different spacing requirements to prevent unwanted crosstalk. For high-index contrast systems (e.g., silicon-on-insulator), relatively small spacings of 1–3 m are typically sufficient.
@@ II-B5 Signal Integrity
Insertion loss is a critical metric for evaluating PIC routing quality, as it directly affects the laser power budget and signal integrity, including signal-to-noise ratio and crosstalk, as illustrated in Fig. 2(f) The primary evaluation metric is the maximum insertion loss along the critical path. Long waveguides and excessive crossings degrade signal integrity and should be minimized wherever possible.
@@ II-C Schematic-Driven PIC Layout
Traditional PIC layout workflows, including manual design and current EPDA tools, are schematic-driven [30]. In this approach, all structures, including crossings and waveguide segments, are instantiated explicitly. Designers must plan routing during the schematic stage and manually insert crossings, while waveguide connections rely on port abutment without explicit physical net definitions.
@@ III LiDAR: Automated PIC Detailed Routing
With an input circuit netlist, as shown in Fig. 3, we present LiDAR, a detailed PIC routing framework based on customized grid-based A∗ search. It efficiently generates curvy waveguide paths with automatic crossing insertion, minimizing maximum insertion loss while ensuring design-rule compliance. The core routing flow has three phases: ➊ Port Access Assignment: assigns ports by orientation and local density to ease routing and reduce congestion; ➋ Iterative Curvy-Aware Routing: routes all nets with a curvy-aware A∗ guided by group-based net ordering; ➌ Crossing Optimization and Refinement: applies local rip-up and reroute to resolve conflicts and optimize crossings, then refines results to remove redundant bends and produce a clean GDS layout.
@@ III-A Accessibility-Enhanced Port Assignment
Port Propagation. Some waveguide ports are inside the bounding box of a device. Since devices are treated as routing blockages, we propagate these internal ports outward to the device boundary based on their orientations, as illustrated in Fig. 4(a). This ensures they are reachable during routing.
Congested Port Spreading. Certain PIC devices feature dense clusters of ports that may overlap on the same routing grid, leading to access conflicts. To alleviate this, we apply a symmetric spreading strategy, where high-density ports are redistributed with a predefined spacing and extension length, as shown in Fig. 4(b). These new access ports are connected to their original locations via sine bends. We add 5 units of extension length for each grid shift, ensuring they occupy distinct routing tracks and satisfying the minimum bend radius.
Staggered Port Access Region Reservation. To prevent waveguides from blocking ports, a region in front of each port is reserved along its orientation (Fig. 4(a)). For multiport devices such as MMIs, assigning the same reserved length to all ports can still cause blockage, while closely spaced parallel waveguides hinder crossing insertion (Fig. 4(c)), since a crossing requires a minimum footprint and cannot be placed consecutively without sufficient spacing. To address this, we introduce the staggered port access region strategy, ensuring adjacent access points are spaced larger than the crossing footprint, leaving sufficient room for accessibility. The reserved length depends on the port’s order among ports of the same orientation, as well as crossing size and bend radius. Ports are grouped by orientation and sorted by coordinates along the perpendicular axis, each assigned an index . The reserved length for port is calculated by:
where is the number of ports in the group. This yields a staggered, mountain-shaped profile that prevents inner ports from being blocked and facilitates escape. If routing fails, the reserved length is adjusted according to the updated routing order in Section IV-B.
@@ III-B Port-Group-based Net Order
III-C Non-Manhattan Waveguide Routing with Curvy-Aware A∗
@@ III-C2 Parametric Curvy-Aware Neighbor Candidate Generation
Here, and represent the horizontal and vertical grid distances required to reach a neighbor at a 90∘ turn. The ceiling function guarantees that the bend radius constraint is met. Note that unlike 45∘ diagonal neighbors in traditional 8-way A∗ search, where is equal to , our neighbors are the endpoint of the 45-degree curves, where and are not equal.
@@ III-C4 Insertion Loss-Aware A∗ Search Cost
LiDAR adopts a customized A∗ search cost function that incorporates insertion loss considerations to improve both routing quality and algorithm efficiency. The total cost of a node in A∗ search is given by the standard formulation: , where denotes the accumulated cost from the source node to the current node , and is the heuristic estimate of the cost from node to the target node . The cost includes an insertion-loss-inspired term , adapted from Eq. (1) with a tunable weight and a group-based congestion penalty (GCP), denoted as to save routing resources for a group of nets. By adjusting the weight parameters, users can flexibly balance between waveguide length and the number of crossings to meet specific design objectives:
@@
Since our method allows non-Manhattan routing, the shortest path is often diagonal. We define a customized heuristic as shown in Eq. (6).
where and represent the minimum and maximum differences in the x and y coordinates between the current node and the target node , respectively. When , we have , and reduces to the standard Manhattan case. When , it indicates that the subsequent path search must contain at least one bend with an angle no smaller than . Therefore, we incorporate the loss of a bend into , which does not overestimate the cost and thus the A∗ router remains admissible, and penalizes the bend and encourages paths that end in orientations suited for port connection.
@@ IV-A Hierarchical Netlist Tree Construction
PIC netlists often exhibit inherent hierarchical structure, where complex systems are composed of repeated functional modules, as shown in Fig. 11. Leveraging this structure is crucial for scalable physical design. To exploit such hierarchy, we construct a Hierarchical Tree by parsing the nested module instantiations defined in the YAML-format netlist, which is similar in style to a SPICE netlist and inherently preserves hierarchical relationships. In this tree, the root node corresponds to the top-level module of the circuit, the leaf nodes represent primitive photonic components, and the internal nodes represent intermediate-level submodules. The resulting tree has arbitrary fanout and depth, depending on the circuit topology. The construction has linear time complexity of , where is the total number of unique modules and components. This hierarchical representation enables global routing planning at multiple levels.
We then perform bottom-up routing for each level of modules progressively. Within each level, the priority of each routing group is determined according to the group priority score defined in Subsection III-B. This multi-level routing approach helps avoid conflicts between different modules, as routing is localized within each submodule before higher-level integration. Similar to standard-cell design in VLSI, the internal routing resources of a module are treated as private and should not be accessed by external nets.
@@ IV-B Routing Conflict Reduction via Spatial-Aware Optimization
Offset Neighbor. Although the algorithm discourages excessive bends via its bend-cost heuristic, it may still add unnecessary bends when ports are slightly misaligned. In tight spaces, such misalignment can cause multiple bends, raising insertion loss and risking routing failure (Fig. 12(a)).
To address this, we add offset neighbors under specific conditions: when the current node is within four bending radii of the target and their orientations differ by 180 degrees. These neighbors correct small misalignments (below the bend radius) between source and target ports. Their locations are computed analytically from the bend radius and grid size, ensuring smooth, fabricable sine-bend waveguides. From an algorithmic perspective, offset neighbors simplify port access, cut redundant detours, and improve search efficiency.
@@
where is a user-defined penalty and is the minimum width required to insert a valid crossing. The penalty is applied over a -wide region around the net segment, discouraging other nets from occupying this reserved space.
@@ IV-C Reduce Routing Conflict via Routing Order Optimization
Although port regions allow staggered access points to reduce conflicts, an improper order may still block target ports with routed waveguides or other regions. For multi-port components, the intra-group routing order is thus crucial. As shown in Fig. 13(a), both orientation and spatial location of ports strongly influence order, as they determine accessibility and potential conflicts.
To address this, we derive the order from the spatial relation between source and target ports (Fig. 13(b)). Let the source port set be with positions . Their mean coordinates are:
Routing order is then based on Manhattan distance from each target port to , with closer ports given higher priority. This method effectively resolves most conflicts, especially in multiport components.
Routing Order Refinement. It is difficult to guarantee an optimal routing order. When nets are routed independently without group information, the method above is inapplicable, making dynamic order adjustment necessary. An improper order may block ports with earlier waveguides, causing escape or access failures. As shown in Fig. 14, routing net1, net2, and net3 independently leads to net2’s failure, as net1’s port region and net3’s waveguide block its path.
To address routing failures, we enhance rip-up and re-route by removing not only nets directly conflicting with the failed path but also those in front of the failed net’s ports to ease congestion. As shown in Fig. 14, we scan along the port orientation, identify conflict order and conflicting nets (starting from ), and use them to adjust routing order. For instance, if net3 is blocked by net2, and net2 by net1, we merge them and uniformly update the routing order from 3 to 1. When net3 and net2 are both blocked by net1 and their order is indistinguishable, we sort them by net names. Although higher-priority nets are routed first, they may still block others. To prevent this, we dynamically extend high-priority target ports. Specifically, the port length is extended by , where is the number of merged nets, the order, and the bend radius, shifting access points and reducing blocking risk. Our extended conflict-resilient hierarchical routing scheme can effectively enhance routability even under tight layout constraints and help accelerate the routing process.
@@ V PIC Routing Benchmark
Benes [38]: Based on Mach-Zehnder interferometers, it adopts a structured, hierarchical Benes topology. It exhibits extremely high crossing density, where a single net may require tens of crossings, challenging both crossing reservation and congestion-aware routing strategies.
Tables II and III summarize the benchmark statistics. Total #CR refers to the number of topological crossings inherent to the circuit and resolved during routing. For Light and GWOR, internal switch crossings are excluded, as we focus solely on the routing between the switch and the I/O interfaces in these benchmarks. Table II (ISPD version) includes spacious, moderately complex layouts, while Table III features more compact benchmarks with higher component density, smaller die sizes, and more crossings, posing greater routing challenges and better reflecting real-world photonic layout constraints. For the Light and GWOR benchmarks, the bending radius, waveguide width, and grid resolution are set to 60 m, 2 m, and 50 m, respectively, to emulate the larger bend radii of the SiN platform. Since these circuits provide abundant routing resources, a larger grid size is adopted to accelerate the search process. For the other benchmarks, these values are set to 5 m, 0.5 m, and 2 m, respectively, reflecting the tighter bends of the SiPh platform. By default, the crossing size is fixed at 10 m 10m.